the equation I = 2.4e-6t defines how current varies with time


1
Expert's answer
2021-02-02T05:03:05-0500
"I=\dfrac{dq}{dt}=2.4e^{-6t}"

The quantity of charge which has to be passed through a cross section of the wire during the time "t_1" to "t_2" is​


"\displaystyle\int_{t_1}^{_2}2.4e^{-6t}dt=[-0.4e^{-6t}]\begin{matrix}
 t_2 \\
 t_1
\end{matrix}"

"=-0.4e^{-6t_2}+0.4e^{-6t_1}"